OpenStudy (anonymous):

solve for x using logarithms(round to 3 decimal places) 5x=100

OpenStudy (anonymous):

i am going to assume that this is \(5^x=100\) in which case use the change of base formula to solve \(b^x=A\) for \(x\) use \(x=\frac{\ln(A)}{\ln(b)}\)

OpenStudy (anonymous):

in your case you want \[x=\frac{\ln(100)}{\ln(5)}\] or you can use \[x=\frac{\log(100)}{\log(5)}\] and then a calculator

OpenStudy (anonymous):

here is what i get, is it right? x=log5 7=1.209